OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 14 –3x + 2y = 1 5 3 y can be any number as there are infinitely many solutions to the system. There is no y value as there is no solution to the system.

OpenStudy (anonymous):

3x + y = 14 -3x + 2y = 1 If you add the two equations together, the x's cancel, leaving you with: 3y = 15 y = 5

OpenStudy (anonymous):

hope i helped

OpenStudy (anonymous):

Thanks so much!